OpenStudy (anonymous):

Granite College has 1,200 female students. This represents 40 percent of the total student body. The number of male students at Granite is

OpenStudy (anonymous):

1,200 female = 40 % x = 100% 1,200 * 100 = 40 * x 12,000 = 4 x x = 3,000 3,000 - female = male 3,000 - 1,200 = 1,800